A person in good physical condition can put out 100 W of useful power for several hours at a stretch, perhaps by pedaling a mechanism that drives an electric generator. Neglecting any problems of generator efficiency and practical considerations such as resting time:

(a) How many people would it take to run a 5.00 kW electric clothes dryer?

(b) How many people would it take to replace a large electric power plant that generates 500 MW?
Answer #1

here,

the power generated by each Person , P = 100 W

a)

for P1 = 5 KW = 5000 W

the number of persons required , N1 = P1 / P

N1 = 5000 /100 peoples

N1 = 50 peoples

b)

for P2 = 500 MW = 5 * 10^8 W

the number of persons required , N2 = P2 / P

N2 = 5 * 10^8 /100 peoples

N2 = 5 * 10^6 peoples